OpenStudy (anonymous):

Determine the slope of the secant line for the curve defined by the equation: f(x)=-7x^2 + 3 for the points determined by x(sub0)=7 and x(sub1)=10.

OpenStudy (anonymous):

at x(sub0)=7 f'(x(sub0))=-7*2*0=0 at x(sub1)=10 f'(x(sub1))=-7*2*1=-14

OpenStudy (anonymous):

for the first on it's -7*2*7=-14*7=-98 and the second it's -7*2*10=-140

OpenStudy (anonymous):

thanks